# Wichitrnithed, August 28, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Solver (DG-SWEM) using CUDA and OpenACC

### v1

@@ 7 Discussion
which is based on some assumptions, e.g., latencies are ignored and computation perfectly overlaps with data transfer [Hager and Wellein (2010)]. In subroutines with low (like most in our code), performance is primarily bounded by data transfer time and will not reach the peak FLOP/s of the GPU. To plot the rooflines, we use the procedure described in [Yang (2020)].
7.1 Wind forcing kernels
As seen in the previous section, a significant performance gain lies in the wind processing kernels inside the met_forcing routine. While they consist of short loops, their performances do vary as seen in Figure 9. Wind kernels 3 and 4 reach reasonable fractions of effective performance bounds, while kernels 1 and 2 exhibit a much lower performance, when only looking at the DRAM / HBM roofline. This disparity can be attributed to the different types of calculation between these two groups of kernels. Kernels 3-5 only do simple arithmetic over each array element, while kernels 1-2 contain nested calls to functions that mostly perform repeated data accesses. This is indicated by a large separation between the HBM and L1/L2 intensities, showing good locality in the higher-level caches. In this case, performance is bound by the L2 bandwidth. As for wind kernel 5, performance is likely limited by the low reported occupancy of only 28% due to the use of many registers.
7.2 Integration kernels
Similar to the wind kernels, integration kernels are bandwidth-bound (Figure 10). However, we do observe good locality in the caches as indicated by the separations. In particular, the volume integration kernel reaches quite close to the L1 bandwidth ceiling, and similarly the L2 ceiling for the edge integration kernel. The main performance inhibitor, however, is likely the low occupancy (only 14% and 20% for edge and volume, respectively) resulting from the high register usage from each kernel (168 and 128 per thread). This is a hard constraint on our code, as we have a large number of variables (which are also likely to grow in the future) required by the various equations in the model. In fact, we had to use the launch_bounds kernel attribute to limit the amount of registers used (255) and increase occupancy, which was originally below 10% for both kernels.
Other memory-related performance factors include non-coalesced (strided) accesses to the state arrays global access patterns. As seen in Table 4, many arrays are arranged with the element id as the second index, which leads to accesses with a stride of 3 (the degrees of freedom) when each element is assigned to a thread. Global access pattern is also an issue for the edge integration kernel, as each edge must access neighboring elements where are not arranged contiguously in memory, a problem inherent in the use of unstructured meshes [Hadade et al. (2020)]. Finally, we also note that using linear () elements is typically insufficient to reach close to peak performance of the GPU [Fuhry (2013), Klöckner et al. (2009), Abdi et al. (2019)].
8 Conclusion and future work
We have created a GPU implementation of DG-SWEM using two different approaches: CUDA Fortran and OpenACC. In both cases, the porting is highly straightforward through the use of modern API on the NVIDIA Fortran compiler. We exploit loop-level parallelism from independent element/edge computations inherent in the DG method, and tested both implementations on the new Grace-Hopper system using realistic flood scenarios. We observe significant speedups over the original CPU-MPI implementation on a single node (144 cores) of 2.3-2.5 for a riverine flooding case, and over 6 for the storm surge cases. We also observe that the OpenACC version, when combined with the CUDA Unified Memory System (using -unified) on Grace-Hopper, requires very few code modifications and can be maintained in the same code base; it also offers a performance close to the current CUDA Fortran version. However, some care is needed in adding specific clauses to remove overhead and prevent a too fine-grained parallelization.
@@ 10 Acknowledgements
This work has been supported by the United States National Science Foundation NSF PREEVENTS Track 2 Program, under NSF Grant Numbers 1855047 This material is also based on work supported by the US Department of Homeland Security under Grant No.2015-ST-061-ND0001-01. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the US Department of Homeland Security.

### v2

@@ 6.2 Performance Evaluation
which makes some assumptions, e.g., latencies are ignored and computation perfectly overlaps with data transfer [Hager and Wellein (2010)]. In kernels with low (like most in our code), performance will, at best, reach only a fraction of the peak FLOP/s of the GPU.
To plot the empirical roofline, we use the procedure described in [Yang (2020)]; this is shown in Figure 5. The AREA, EDGE, and GATHER kernels achieve compute throughputs of 1.9 TFLOPs/s, 1.3 TFLOPs/s, and 0.24 TFLOPs/s, respectively. Other relevant performance characteristics from Nsight Compute are shown in Table 6. A few observations can be made. First, the INT instruction count shows that a large portion of the computation are in fact array index calculations. While this is largely unavoidable, they do lower the measured FLOP rate, as we cannot assume they overlap with FP32 instructions. Second, the achieved occupancy for the first two kernels are quite low due to the high register usage, and this can also limit performance. While it is possible to tune register usage through the -maxregcount compiler flag, we found that the penalty from spilling largely offsets the speedup from higher occupancy. We also note that using linear () elements is typically insufficient to reach close to peak performance of the GPU [Fuhry (2013), Klöckner et al. (2009), Abdi et al. (2019)].
| Kernel | Registers per thread | Occupancy (%) | FP32 Inst. | INT Instr. |
| EDGE | 249 | 9.4 | 197 | |
| AREA | 163 | 14.9 | ||
| GATHER | 72 | 40.6 |
6.3 Lines of code changed
To get a rough estimate of the level of programming effort, we compare the total lines of code before and after incorporating OpenACC (we ignore blank lines). Lines of code increase by 147, from 48,674 to 48,821; this change includes the addition OpenACC directives and other modifications such as the flux-gathering routine and MPI packing/unpacking. Table 7 shows the breakdown of the specific OpenACC directives used.
| Directive | Description | # of lines |
!$acc parallel |
Loop parallelism | 29 |
!$acc routine |
Convert of CPU to GPU routines | 15 |
!$acc loop seq |
Serialize loops | 27 |
!$acc atomic |
Atomic updates | 15 |
!$acc wait |
Barrier | 8 |
| Continuation lines | 4 | |
| Total | 98 |
7 Conclusion
We have ported DG-SWEM to NVIDIA GPUs using OpenACC Fortran in a basic manner by exploiting the data parallelism inherent in the explicit-time DG method through element and edge computations . We tested the code on multiple nodes on the NVIDIA Grace Hopper system using a large flooding test case and observed significant speedups over the original CPU implementation. When combined with the CUDA Unified Memory System, the porting requires very few code modifications and can be maintained in the same code base; this is a marked improvement from our previous porting framework [DuChene et al. (2011)]. Finally, we note that there are potential performance issues that can be further investigated on top of this basic parallelization. First, a more detailed benchmark on multi-node scaling can be done to observe the code’s relative performance, and whether further optimizations like overlapping computation and communication will be beneficial. Second, we can explore alternative ways to structure edge integration to exploit GPU shared memory and improve data locality among neighboring edges.
@@ 9 Acknowledgements
This work has been supported by the United States National Science Foundation NSF PREEVENTS Track 2 Program, under NSF Grant Numbers 1855047. This material is also based on work supported by the US Department of Homeland Security under Grant No.2015-ST-061-ND0001-01. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the US Department of Homeland Security.